OpenStudy (anonymous):

One triangle on a graph has a vertical side of 7 and a horizontal side of 12. Another triangle on a graph has a vertical side of 28 and a horizontal side of 48. Could the hypotenuses of these two triangles lie along the same line? Yes, because they are similar triangles Yes, because all triangles can fit along this line No, because they need to be the same size No, because they are not similar triangles

OpenStudy (anonymous):

i think yes, but idk between a or b

OpenStudy (michele_laino):

we can speak about a similitude ratio, since the ratio between corresponding sides is: \[\frac{{48}}{{12}} = \frac{{28}}{7} = ...?\]

OpenStudy (michele_laino):

more precisely such ratio holds also between the corresponding hypotenuses: \[\large \begin{gathered} hypotenus{e_1} = \sqrt {{7^2} + {{12}^2}} = \sqrt {49 + 144} \hfill \\ \hfill \\ \hfill \\ hypotenus{e_2} = \sqrt {{{\left( {4 \times 7} \right)}^2} + {{\left( {4 \times 12} \right)}^2}} = \sqrt {{4^2} \times \left( {{7^2} + {{12}^2}} \right)} = \hfill \\ \hfill \\ = 4\sqrt {{7^2} + {{12}^2}} = 4 \times hypotenus{e_1} \hfill \\ \end{gathered} \] so the two triangles are similar each other